OpenStudy (anonymous):

Solve the equation or formula for the indicated variable. S=5r^2t for t A. t=s/5-r B. t=25r/s C. t=r^2-5s D. t=s/5r^2

OpenStudy (anonymous):

hmmm none of these if it is really \( S=5r^t \)

OpenStudy (lgbasallote):

actually i think there is...

OpenStudy (anonymous):

sorry! its S-5r^2t

OpenStudy (anonymous):

hehe I knew it

OpenStudy (anonymous):

oh maybe it is \(S=5r^2t\)

OpenStudy (anonymous):

S= i mean

OpenStudy (anonymous):

ask Emily to write it again

OpenStudy (anonymous):

S=5r^2t

OpenStudy (anonymous):

divide both sides by \(5r^2\) to get \(\frac{S}{5r^2}=t\)

OpenStudy (anonymous):

what if it were 2t?

OpenStudy (anonymous):

ahhhh.. Thanks satellite73

OpenStudy (anonymous):

if it were \(S=5r^{2t}\) you get \(t=\frac{\ln(\frac{S}{5})}{2\ln(r)}\)

OpenStudy (anonymous):

which is why it seemed unlikely yw ej

OpenStudy (anonymous):

I just want to make sure that it isn't what is being sought for...